OpenStudy (anonymous):

How would you put the below system into y= ? form? x = -(3/2)y x = (3/2)y - 2

OpenStudy (anonymous):

divide each side by -(3/2) add 2 to each side, then divide each side by (3/2)

OpenStudy (anonymous):

so y=x/-(3/2) and y=(x-2)/(3/2)

OpenStudy (anonymous):

nono (x+2)/(3/2)

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

as they are both = x -(3/2)y = (3/2)y - 2 divide by -(3/2): y = -y -2/-(3/2) 2y = 4/3 y = 2/3